OpenStudy (anonymous):

An initial investment of $50,692 earns 6% interest compounded monthly. How many months will it take for the investment to double? Show work

OpenStudy (anonymous):

A formula for calculating compound interest is A = P(1+r/n)^n Where, A = final amount P = principal amount (initial investment) r = annual nominal interest rate (as a decimal) (it should not be in percentage) n = number of times the interest is compounded per year t = number of years This is the starting point.

OpenStudy (anonymous):

A = P(1+r/n)^nt Sorry, missed a t

OpenStudy (anonymous):

so how do u solve this

OpenStudy (anonymous):

?

OpenStudy (anonymous):

So, you know some information already. Double the initial investment is 101384, 101384 = 50692 (1 + 0.06/12)^nt

OpenStudy (anonymous):

Do a little calculating and rearranging, you get: 101384/50692 = (1.005)^nt 2 = (1.005)^nt Should be looking a little easier by now.

OpenStudy (anonymous):

The last step is to find out how many months (nt) it takes. You can do this as follows: ln(2) = ln((1.005)^nt) ln(2) = ln(1.005) * nt ln(2)/ln(1.005) = nt You may need a calculator for this, but is should work out to around 139 months.